The algebra test had a mean of 82 and a standard deviation of 8. What percentage of scores fall between 82 and 90.
Answer #1

Solution :

Given that ,

mean = \mu = 82

standard deviation = \sigma = 8

P( 82 < x <90) = P[(82 - 82) / 8 ) < (x - \mu) /\sigma  < (90 - 82) / 8) ]

= P( 0 < z < 1)

= P(z < 1 ) - P(z < 0)

Using z table,

= 0.8413 - 0.5

= 0.3413